OpenStudy (anonymous):

Euler discovered that the polynomial p(n) = n^2 - n + 41 yields prime numbers for many small positive integer values of n. What is the smallest positive integer n for which p(n) and p(n+1) share a common factor greater than 1?

OpenStudy (freckles):

@ganeshie8 you might like this question

OpenStudy (freckles):

I don't know if we can do anything with this... but I have: \[\frac{n+1}{2} p(n)-\frac{n-1}{2}p(n+1)=41\]

OpenStudy (freckles):

Did a couple of polynomial division problems to get that... You know like euclidean algorithm stuff.

ganeshie8 (ganeshie8):

n^2 -n + 41 = n(n-1) + 41 plugin n = 41 or 42 :P

ganeshie8 (ganeshie8):

i think this is @cwrw238 's favorite problem

ganeshie8 (ganeshie8):

wait sry i misread the question, it looks a bit more involved..

OpenStudy (freckles):

I think from that equation above n is odd just thinking I want the (n+1)/2 and (n-1)/2 to be integers

OpenStudy (freckles):

But also like I said before I don't know if that equation helps yet

OpenStudy (freckles):

But also like I said before I don't know if that equation helps yet

OpenStudy (freckles):

for what @ganeshie8 said n=41 does give gcd 41 but I think for n=42 it gives gcd 1 I wonder if n=41 is the lowest positive integer n such that the gcd is >1

ganeshie8 (ganeshie8):

Ahh that actually works! p(41) and p(42) are actually composite with a gcd of 41

ganeshie8 (ganeshie8):

p(n) = `n(n-1)` + 41 plugging in n=41 or 42 makes the first term have a factor of 41

OpenStudy (freckles):

ah p(41)=41(40+1) p(42)=41(42+1)

OpenStudy (freckles):

@danyboy9169 just mentioning you here because I think you opened up a new question. just wanted to make sure you see the discussion here.

OpenStudy (anonymous):

Yes, I did bumped it again as this is Number Theory descrete math, been searching for last 4 hrs now.. but if I know that for two consecutive nos. gcd is always 1.

OpenStudy (freckles):

\[\frac{n+1}{2} p(n)-\frac{n-1}{2}p(n+1)=41 \\ \text{ this means we can either have } \gcd(p(n),p(n+1))=1 \\ \text{ or } \gcd(p(n),p(n+1))=41\] those are the only two possibilities ------------ p(41) and p(42) are not consective though 41 and 42 are though

ganeshie8 (ganeshie8):

Oh you're using that linear combination thingy here ax + by = c for above equation to make sense, we must have gcd(a,b) | c only 1 and 41 divide 41 so we're done.

OpenStudy (freckles):

ci (yes; I think ci is spanish word for yes (it might be si))

OpenStudy (anonymous):

thanks a ton freckles and ganeshie8

ganeshie8 (ganeshie8):

as i see now thats actually pretty clever !